OpenStudy (anonymous):

Use the substitution method to solve the linear system given. Express the answer as an ordered pair of the form (x,y) x + y = 5 2x + y = 6

OpenStudy (campbell_st):

rewrite the 1st equation let y = 5 - x and substitute in to the 2nd equation 2x + ( 5 -x) = 6 then 2x + 5 - x = 6 x + 5 = 6 subtract 5 from both sides x = 1 substitute x = 1 into the 1st equation 1 + y = 5 then y = 4 so the ordered pair is (1, 4)